A firm has a current capital structure consisting of $400,000 of 6 percent annual interest debt...

  1. A firm has a current capital structure consisting of $400,000 of 6 percent annual interest debt and 50,000 shares of common stock. The firm's tax rate is 21 percent on ordinary income. If the EBIT is expected to be $200,000, two EBIT-EPS coordinates for the firm's existing capital structure are ________.

    ($0, $24,000) and ($200,000, $1.82)

    ($24,000, $0) and ($200,000, $3.52)

    ($24,000, $0) and ($200,000, $2.78)

    ($24,000, $0) and ($200,000, $0.74)
Answer #1

EBIT=24000
EPS=(24000-400000*6%)*(1-21%)/50000=0

EBIT=200000
EPS=(200000-400000*6%)*(1-21%)/50000=2.7808

($24,000, $0) and ($200,000, $2.78)